OpenStudy (goformit100):

At a local supermarket, a box of cereal usually costs 5.56 dollars. This week, the supermarket sells the box of cereal for 3.42 dollars. How much money will you save if you buy this cereal at this supermarket?

OpenStudy (anonymous):

5.56-3.42=2.14

OpenStudy (anonymous):

5.56 - 3.42 = 2.14

OpenStudy (anonymous):

Yeah, what he said lol

OpenStudy (yrelhan4):

Medal goes to juan this time. :P

OpenStudy (anonymous):

xc mkay

OpenStudy (goformit100):

its not a medal game.

OpenStudy (anonymous):

Get a calculator. 5.56 - 3.42 = 2.14 Don't forget the $ sign or you will lose marks.

OpenStudy (goformit100):

Please Help me ... I am unable to understand the question

OpenStudy (goformit100):

i don't prefer calculator

OpenStudy (anonymous):

5.56 - 3.42 = (6 - 2 = 4) (5-4 = 1) (5 - 3 = 2) Decimal point stays the same 2.14

OpenStudy (anonymous):

HOW do you give medals?

OpenStudy (anonymous):

You click on the medal button on the right corner of the reply

OpenStudy (yrelhan4):

Best response button*

OpenStudy (yrelhan4):

Bring on the medals. \m/ :P

OpenStudy (anonymous):

Whatever. :P